For a vehicle that can accelerate from 0-60 mph in 6.2 seconds, how far would it need to travel to go from 0 mph to 58 mph?
Answer by Alan3354(30993)

For a vehicle that can accelerate from 0-60 mph in 6.2 seconds, how far would it need to travel to go from 0 mph to 58 mph?
-----------------------------
60 mph = 88 ft/sec
velocity = acceleration * time
v = at
a = v/t = 88/6.2 = 440/31 ft/sec/sec
-------------------
58 mph = 58*22/15 ft/sec = 1276/15 ft/sec
--------------------
t = v/a = (1276/15)/(440/31) = 899/150 seconds
--------------
distance = (acc * time^2)/2
s = (1/2)at^2
s = (1/2)*(440/31)*(899/150)^2
s = 254.9 feet